OpenStudy (anonymous):

Double-check please: The sail of a boat is in the shape of a right triangle. ----------------------------------- Which expression shows the height, in meters, of the sail?

OpenStudy (anonymous):

Pythagurus

OpenStudy (anonymous):

My answer: 8/cosec 35deg

OpenStudy (anonymous):

lol he might have been a guru, but it was Pythagoras

OpenStudy (shubhamsrg):

why cosec?? it should be "cos"

OpenStudy (anonymous):

\[\cos(35)=\frac{8}{?}\] \[?=\frac{8}{\cos(35)}\]

OpenStudy (shubhamsrg):

i mean either 8cos 35 or 8/sec 35

OpenStudy (anonymous):

I don't know 1 more length so I can't use pythagoras's theorem. And no-one of the answers have Cos, even though i concluded to that as-well.

OpenStudy (anonymous):

then a calculator

OpenStudy (shubhamsrg):

if unknown side is x then cos 35 = x/8,,and not 8/x

OpenStudy (anonymous):

If Cos isn't an applicable answer it must be something else then, no?

OpenStudy (shubhamsrg):

8/(sec 35)

OpenStudy (shubhamsrg):

its the same as 8 cos 35

OpenStudy (anonymous):

howso?

OpenStudy (shubhamsrg):

1/cos x =sec x ..this is basic trigonometry sir..

OpenStudy (anonymous):

You can blame that on the poor Geometry teacher I have at the moment :p

OpenStudy (shubhamsrg):

you mean trigonometry teacher? :P

OpenStudy (anonymous):

Nope. They teach students trig functions in a Geo class. Thats America for ya.

OpenStudy (shubhamsrg):

hmmn,,but still the teacher cant be that dumb,, anyways,,it was nice to help.. ^_^

OpenStudy (anonymous):

If you had her you'd understand xD thank you!

OpenStudy (shubhamsrg):

lol XD